OpenStudy (anonymous):

Hi I really need someone's help I've been struggling with this question for what seems like forever: Celina is buying fabric scraps to make koozies for her glasses. Each glass is cylindrical with a diameter of 4 inches and a height of 8 inches. The koozies will cover only the sides of the glasses. What is the amount of fabric she will use? What does this amount represent? Use for . Show your work.

OpenStudy (mathstudent55):

Each glass is shaped like a cylinder. You need to find the lateral area of a cylinder, and multiply it by the number of glasses. The lateral surface of a cylinder is shaped like a rectangle that has the circumference of the base as one side and the height of the cylinder as the other side.

OpenStudy (mathstudent55):

Think of a cylindrical can with a label all around it. If you cut the labal vertically, along the height of the cylinder, and you unwind it and flatten it, you'll have a rectangle that has C (the circumference of the base) and h (the height of the cylinder) as its length and width. 

OpenStudy (anonymous):

could you explain further?

OpenStudy (mathstudent55):

Your problem states to use [blank] for [blank]. I assume it is asking you to use something for pi, but you have blanks instead. I am using 3.14 for pi. The area of the side of each glass is C * h = (pi)d * h = (3.14)(4 in.)(8 in.) = 100.48 in.^2 The area of each koozie is 100.48 in.^2 This amount represents the lateral area of each glass. In other words, it's the surface area of a cylinder leaving out the top and bottom bases.